OpenStudy (anonymous):

Which of the following is the radical expression of 2d^7/10?

hartnn (hartnn):

what u got ?

OpenStudy (anonymous):

I'm working on it :)

OpenStudy (anonymous):

by the way, hello @hartnn !!!!!!

hartnn (hartnn):

\(\huge a^{\frac{m}{n}}=^n\sqrt{a^m}\) hi :)

OpenStudy (anonymous):

C

OpenStudy (anonymous):

That makes more sense to you than us... you didn't post the choices yet ;)

OpenStudy (anonymous):

My error ;)

OpenStudy (anonymous):

7√2d^10 10√2d^7 2^10√d^7 2^7√d^10

hartnn (hartnn):

\(\huge (2d)^{7/10}\) was this the question ?

OpenStudy (anonymous):

The 2d was not in parentheses :)

hartnn (hartnn):

\(\huge 2(d)^{7/10}=2\sqrt[10]{d^7}\) now this seems right.

OpenStudy (anonymous):

I GOT IT RIGHT?!!!!!!!

hartnn (hartnn):

yup

OpenStudy (anonymous):

THANK YOU

hartnn (hartnn):

welcome :)

hartnn (hartnn):

u can learn how to write exponents and square roots, using latex....just to make questions clearer....